jack is three times as old as Jill the total of their ages is 72. Find the age of each person
Answer by mananth(16946)

jill ----------- x years
Jack = 3x
Jack's age + Jill's age =72
x+3x=72
4x=72
/4
x=72/4
x=18 Jill's age
Jack's age = 3x=3*18= 54 years